# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v2

@@ 3.3 Enforcing physics consistency
Let $\mathbf{u}:\mathbb{R}^{3}\!\to\!\mathbb{R}^{3}$ be the velocity field and denote its value at any point $\mathbf{x}=(x_{1},x_{2},x_{3})$ by $\mathbf{u}(\mathbf{x})$. The local rotation of the velocity field is characterized by the vorticity, which is defined as the curl of the velocity:
$\bm{\omega}(\mathbf{x})=\nabla\times\mathbf{u}(\mathbf{x})=\begin{pmatrix}\frac{\partial u_{3}}{\partial x_{2}}-\frac{\partial u_{2}}{\partial x_{3}}\\ \frac{\partial u_{1}}{\partial x_{3}}-\frac{\partial u_{3}}{\partial x_{1}}\\ \frac{\partial u_{2}}{\partial x_{1}}-\frac{\partial u_{1}}{\partial x_{2}}\end{pmatrix}.$ (10)
A well-known mathematical property of the vorticity (and more generally the curl of a field) is that its divergence vanishes, i. e., (×$\nabla\cdot(\nabla\times\mathbf{u})=0$. Intuitively, this means that vorticity has no monopole-like sources or sinks: vortex lines can only form closed loops or extend to infinity, never beginning or ending at a point.
Models that directly predict the vorticity field $\bm{\hat{\omega}}(\mathbf{x})$ without further structure cannot guarantee $\nabla\cdot\bm{\hat{\omega}}=0$, and thus lead to unphysical predictions that introduce spurious sources and sinks of vorticity. PINN typically address this by incorporating the residuals of the PDE and boundary conditions into the loss function as soft constraints. By contrast, we enforce the divergence-free condition as an architectural hard constraint by explicitly parameterizing the vorticity field as $\bm{\hat{\omega}}(\mathbf{x})=\nabla\times\mathbf{\hat{u}}(\mathbf{x})$, where $\mathbf{\hat{u}}$ is the modeled velocity field. In other words, we define the vorticity as the curl of the modeled velocity field, which guarantees $\nabla\cdot\bm{\hat{\omega}}=0$ by construction.
@@
$\frac{\partial\hat{u}_{i}}{\partial x_{j}}(\mathbf{x},\,\cdot)$ $\approx\frac{\hat{u}_{i}(\mathbf{x}+\delta\,\mathbf{e}_{j},\,\cdot)-\hat{u}_{i}(\mathbf{x}-\delta\,\mathbf{e}_{j},\,\cdot)}{2\,\delta},$ (11)
where $\hat{u}_{i}$ is the i-th (scalar) velocity output, $\mathbf{e}_{j}$ is the unit vector in the $x_{j}$-direction, and $\delta$ is the step-size. Here, the dot in $\mathbf{\hat{u}}(\mathbf{x},\,\cdot)$ indicates that the prediction also depends on additional inputs such as the anchor token’s key and value vectors, which we omit for notational simplicity. To approximate the partial derivatives w. r. t. each of the three spatial coordinates, we perform six additional forward passes, each applying an offset $+\delta\mathbf{e}_{j}$ and $-\delta\mathbf{e}_{j}$ to the input position.
@@
In order to effectively train neural networks, input and target data is usually shifted and scaled into some appropriate range. Since the vorticity is defined through spatial derivatives of velocities w. r. t. the input coordinates in the raw physics space, the effect of scaling must be taken into account. This is because we compute the spatial derivatives of the neural network, i. e. in network space. Hence, when scaling the 3D input positions by a vector $\mathbf{a}$ (from physics to network space) and the outputs by vector $\mathbf{b}$ (from network to physics space), then the $3\times 3$ network-space Jacobian matrix of partial derivatives must be multiplied by the outer-product matrix $\mathbf{b}\mathbf{a}^{T}$, resulting in the physics-space Jacobian.
@@ 4 Experiments
We begin our experiments by showing that neural surrogates do not need to process the full mesh to obtain accurate models (Section 4.2), which enables efficient training on large problem-scales. Using this insight, we design Anchored-Branched Universal Physics Transformers (AB-UPT) by adapting the UPT architecture (Alkin et al., 2024a) to handle complicated physical dynamics of CFD simulations with million-scale meshes (Section 4.3). To demonstrate the effectiveness of AB-UPT, we compare it against other neural surrogate models in Section 4.4. Finally, we demonstrate how AB-UPT can tackle practical use-cases such as calculating aerodynamic drag and lift coefficients (Section 4.5), simulating directly from a CAD geometry without the expensive creation of a simulation mesh (Section 4.6) as well as enforcing physical consistency via a divergence-free vorticity formulation (Section 4.7). Additional experiments are provided in Appendix A. The experimental setup for AB-UPT is outlined in Appendix B.3
4.1 Datasets
We consider three computational fluid dynamics datasets for automotive aerodynamics: ShapeNetCar (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b). Table 2 outlines properties of these datasets.
@@ 4.2 Efficient training on million-scale CFD meshes
Figure 5: Train and test losses with varying number of training inputs. Contrary to numerical solvers, neural surrogates do not require million-scale meshes to train accurate models. Training on more points beyond a dataset-specific threshold results in overfitting and requires larger datasets (either in terms of data-scale, i.e., number of simulations or problem-scale, i.e., complexity of the problem) to benefit from the increased modeling capacity when using more points, which is costly. This can be observed in the first and second row where AhmedML and DrivAerML have the same data scale (400 train simulations) but AhmedML has smaller problem scale, resulting in overfitting. Note that increasing data- or model-scale is orthogonal to our work. Although runtime increases with a larger number of inputs, the amount of GPU-hours required is still low, and scaling to even larger problem-scales could be easily implemented via sequence parallelism (Liu et al., 2023), which allows scaling quadratic self-attention to millions of tokens.
@@
While such approaches have been explored in the past (Alkin et al., 2024a; b), many neural surrogate architectures focus on linear attention mechanisms (Li et al., 2023a; Wu et al., 2024) or computation parallelization (Luo et al., 2025) to be able to train with millions of inputs. However, such approaches would still be prohibitively expensive. For example, extrapolating the reported GPU requirements for training Transolver++ (Luo et al., 2025) to the 140 million inputs poings from DrivAerML (Ashton et al., 2024b) would necessitate roughly 256 GPUs.
We show in Figure 5 that it is feasible to train standard transformer self-attention to accurately predict simulation results on million-scale meshes by training on lower resolutions, outperforming linear attention architectures (such as Transolver). As self-attention requires quadratically more compute w. r. t. the number of inputs, computational demands increase. However, the number of GPU-hours required to train models is still low (models with 16 thousand surface/volume points train in less than 10 GPU-hours). Given even more complicated physical simulations, it would be no issue to scale our AB-UPT to more GPUs via sequence parallelism (Liu et al., 2023), which distributes tokens among multiple GPUs, to allow processing millions of tokens despite the quadratic complexity of self-attention. Due to the low amount of GPU-hours required, even for physical dynamics that arise from meshes with over 100 million cells, there would be no issue with scaling our model to even more complicated problems. For reference, foundation models in domains such as computer vision (e. g., Oquab et al. (2023)) use 10 to 100 thousand GPU-hours, and compute requirements for language models can be millions of GPU-hours (e. g., Dubey et al. (2024)).
Note that random subsampling as we apply to create a low-resolution input mesh, is also a form of data augmentation where a network will never see the same constellation of inputs twice during training. As high-quality numerical CFD simulations are expensive to obtain (e. g., DrivAerML costs \$1M, see Table 2), datasets often have less than a thousand samples, where this form of regularization helps to prevent overfitting.
@@ 4.3 Model design
We design AB-UPT, a successor of UPT that preserves useful properties of UPT, such as using a neural field for decoding and being mesh independent (i.e., the input is decoupled from the output). We observe (in Table 3) that UPT is not able to model complicated dynamics such as the vorticity well, and gradually change the UPT architecture to employ techniques that allow efficient scaling to large problem sizes and obtain significantly higher accuracies than a vanilla UPT. We visualize design choices in Table 3 and motivate the design decisions below.
Table 3: Impact of design steps for transforming UPT into AB-UPT. Performance is measured as relative L2 error (in %) for surface pressure $\mathbf{p}_{s}$, volume velocity $\mathbf{u}$ and volume vorticity $\bm{\omega}$. Train speed measures seconds per 100 update steps with batch size 1. Test speed measures seconds to create a prediction for the full simulation mesh. Speeds are measured on a NVIDIA H100 GPU. The study is conducted on AhmedML.
@@
$\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$
baseline UPT 4.51 3.26 51.06 7.0 0.2
macro + large decoder 3.90 2.62 13.69 14.7 1.4
design + decoder-only with self-attn 3.42 2.15 6.67 14.9 163
re-add + anchor attention 3.42 2.15 6.67 14.9 2.6
properties + decouple geometry encoding 3.20 2.01 6.47 16.1 2.5
micro + split surface/volume branch 3.50 2.08 6.62 11.0 1.5
design + cross-branch interactions 3.24 2.00 6.69 11.0 1.5
+ branch-specific decoders 3.14 1.93 6.46 11.0 1.5
@@
Starting with a UPT model, we train it on AhmedML (Ashton et al., 2024a) and evaluate surface pressure $\mathbf{p}_{s}$, volume velocity $\mathbf{u}$, and volume vorticity $\bm{\omega}$. UPT struggles with complex decoding tasks (e. g., decoding vorticity) as the decoder consists of a single cross-attention block, and hence, the vast majority of the compute is allocated to encoding.
@@
Subsequently, we aim to improve speed and accuracy. For automotive CFD simulations, we aim to predict surface and volume variables where we use an equal amount of surface and volume tokens. However, these two modalities differ in the way they are calculated in a classical simulation where surface variables have boundary condition $\mathbf{u}=0$. Therefore, we split surface and volume positions into separate branches (“+ split surface/volume branch”). This halves the number of tokens of each individual branch, resulting in faster training/inference speed and decreased performance, particularly for surface variables.
@@ 4.4 Benchmarking AB-UPT against other neural surrogate models
Table 4: Relative L2 errors (in %) of surface pressure $\mathbf{p}_{s}$, volume velocity $\mathbf{u}$ and volume vorticity $\bm{\omega}$. For each model, we indicate if the model has the neural field property, and whether they are mesh independent. Lower percentage values indicate better performance in terms of L2 error. We provide the results for ShapeNetCar, AhmedML, and DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin.
ShapeNetCar AhmedML DrivAerML Neural field Mesh independent
$\mathbf{p}_{s}$ $\mathbf{u}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\bm{\omega}$
PointNet 12.09 3.05 8.51 5.43 67.30 23.63 28.12 880.9
GRAPH U-NET 10.33 2.49 6.42 4.14 53.53 16.13 17.98 380.4
GINO 13.28 2.53 7.90 6.23 71.81 13.03 40.58 98.5
LNO 9.05 2.29 12.91 7.56 72.05 20.51 23.27 173.0
UPT 5.60 1.40 3.90 2.62 13.69 7.85 9.02 189.6
OFormer 7.05 1.61 3.55 2.21 7.78 4.48 6.64 60.4
Transolver 6.50 1.53 3.66 2.19 8.12 4.81 6.78 38.4
Transformer 5.01 1.24 3.41 2.09 6.76 4.35 6.21 47.9
AB-UPT 4.81 1.18 3.00 1.88 6.45 3.82 5.93 35.1
@@
In this section, we compare the performance of AB-UPT against established neural-surrogate models: Table 4 presents the relative L2 error for surface pressure, volume velocity, and vorticity. We compare AB-UPT against the following models: PointNet (Qi et al., 2017a), Graph U-Net (Gao & Ji, 2019), GINO (Li et al., 2023b), LNO (Wang & Wang, 2024), UPT (Alkin et al., 2024a), OFormer (Li et al., 2023a), Transolver (Wu et al., 2024), Transformer (Vaswani et al., 2017). We do not consider Transolver++ as a baseline due to reproducibility issues, which we explain in Appendix B.4. These comparisons are conducted across three diverse datasets: ShapeNetCar (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b). The full experimental setup is detailed in Appendix B. Table 10 in Appendix B presents the training hyperparameters and the number of model parameters for each model. Additionally, Appendix C provides a statement on the fairness of our experimental setup in comparison to (Wu et al., 2024). Comprehensive results, including wall shear stress and volume pressure, are reported in Table 7 in Appendix A.1.
@@
Table 4 shows that AB-UPT consistently outperforms all other neural surrogate models across every evaluated metric and dataset, often by a substantial margin. Moreover, we observe that most baselines struggle with properly modeling the complex vorticity. AB-UPT, however, maintains strong performance for both surface and volume metrics, highlighting its robustness. For most metrics, either a plain Transformer or Transolver is the second-best-performing model. However, due to the full self-attention mechanism of the Transformer, this model will not be able to scale to industry-standard simulation meshes. Transolver, on the other hand, would require sequence parallelism over multiple GPUs to handle large simulation meshes (as has been explored in (Luo et al., 2025). We can conclude that AB-UPT is the best performing model for neural-surrogate automotive CFD simulation, while also having favorable properties concerning problem scaling.

### v3

@@ 3.3 Enforcing physics consistency
Let $\bm{u}:\mathbb{R}^{3}\!\to\!\mathbb{R}^{3}$ be the velocity field and denote its value at any point $\bm{x}=(x_{1},x_{2},x_{3})$ by $\bm{u}(\bm{x})$. The local rotation of the velocity field is characterized by the vorticity, which is defined as the curl of the velocity:
$\bm{\omega}(\bm{x})=\nabla\times\bm{u}(\bm{x})=\begin{pmatrix}\frac{\partial u_{3}}{\partial x_{2}}-\frac{\partial u_{2}}{\partial x_{3}}\\ \frac{\partial u_{1}}{\partial x_{3}}-\frac{\partial u_{3}}{\partial x_{1}}\\ \frac{\partial u_{2}}{\partial x_{1}}-\frac{\partial u_{1}}{\partial x_{2}}\end{pmatrix}.$ (10)
A well-known mathematical property of the vorticity (and more generally the curl of a field) is that its divergence vanishes, i. e., (×$\nabla\cdot(\nabla\times\bm{u})=0$. Intuitively, this means that vorticity has no monopole-like sources or sinks: vortex lines can only form closed loops or extend to infinity, never beginning or ending at a point.
Models that directly predict the vorticity field $\bm{\hat{\omega}}(\bm{x})$ without further structure cannot guarantee $\nabla\cdot\bm{\hat{\omega}}=0$, and thus lead to “unphysical” predictions that introduce spurious sources and sinks of vorticity. Physics-Informed Neural Networks (PINN) typically address this by incorporating the residuals of the PDE and boundary conditions into the loss function as soft constraints. By contrast, in this work we show that we can enforce the divergence-free condition as an architectural hard constraint by explicitly parameterizing the vorticity field as $\bm{\hat{\omega}}(\bm{x})=\nabla\times\hat{\bm{u}}(\bm{x})$, where $\hat{\bm{u}}$ is the modeled velocity field. In other words, we define the vorticity as the curl of the modeled velocity field, which guarantees $\nabla\cdot\bm{\hat{\omega}}=0$ by construction.
@@
$\frac{\partial\hat{u}_{i}}{\partial x_{j}}(\bm{x},\,\cdot)$ $\approx\frac{\hat{u}_{i}(\bm{x}+\delta\,\bm{e}_{j},\,\cdot)-\hat{u}_{i}(\bm{x}-\delta\,\bm{e}_{j},\,\cdot)}{2\,\delta},$ (11)
where $\hat{u}_{i}$ is the i-th (scalar) velocity output, $\bm{e}_{j}$ is the unit vector in the $x_{j}$-direction, and $\delta$ is the step-size. Here, the dot in $\hat{\bm{u}}(\bm{x},\,\cdot)$ indicates that the prediction also depends on additional inputs such as the anchor token’s key and value vectors, which we omit for notational simplicity. To approximate the partial derivatives w. r. t. each of the three spatial coordinates, we perform six additional forward passes, each applying an offset $+\delta\bm{e}_{j}$ and $-\delta\bm{e}_{j}$ to the input position.
@@
In order to effectively train neural networks, input and target data is usually shifted and scaled into some appropriate range. Since the vorticity is defined through spatial derivatives of velocities w. r. t. the input coordinates in the raw physics space, the effect of scaling must be taken into account. This is because we compute the spatial derivatives of the neural network, i. e. in network space. Hence, when scaling the 3D input positions by a vector $\bm{a}$ (from physics to network space) and the outputs by vector $\bm{b}$ (from network to physics space), then the $3\times 3$ network-space Jacobian matrix of partial derivatives must be multiplied by the outer-product matrix $\bm{b}\bm{a}^{T}$, resulting in the physics-space Jacobian.
@@ 4 Experiments
We begin our experiments by showing that neural surrogates do not need to process the full mesh to obtain accurate models (Section 4.2), which enables efficient training on large problem-scales. Using this insight, we design Anchored-Branched Universal Physics Transformers (AB-UPT) by adapting the UPT architecture (Alkin et al., 2024a) to handle complicated physical dynamics of CFD simulations with million-scale meshes (Section 4.3). To demonstrate the effectiveness of AB-UPT, we compare it against other neural surrogate models in Section 4.4. Next, we demonstrate how AB-UPT can tackle practical use-cases such as calculating aerodynamic drag and lift coefficients (Section 4.5), simulating directly from a CAD geometry without the expensive creation of a simulation mesh (Section 4.6) as well as enforcing physical consistency via a divergence-free vorticity formulation (Section 4.7). Finally, we show scaling behavior of AB-UPT w.r.t. problem, model and data scales in Section 4.8. Additional experiments are provided in Appendix A. The experimental setup for AB-UPT is outlined in Appendix B.3
@@
4.1 Datasets
We consider three computational fluid dynamics datasets for automotive aerodynamics: ShapeNet-Car (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b). Table 2 outlines properties of these datasets.
@@ 4.2 Efficient training on million-scale CFD meshes
Figure 5: Train and test losses with varying number of training inputs. Contrary to numerical solvers, neural surrogates do not require million-scale meshes to train accurate models. Training on more points beyond a dataset-specific threshold results in overfitting and requires larger datasets (either in terms of data-scale, i.e., number of simulations or problem-scale, i.e., complexity of the problem) to benefit from the increased modeling capacity when using more points, which is costly. This can be observed in the first and second row where AhmedML and DrivAerML have the same data scale (400 train simulations) but AhmedML has smaller problem scale, resulting in overfitting. Note that increasing data- or model-scale is orthogonal to our work. Although runtime increases with a larger number of inputs, the amount of GPU-hours required is still low, and scaling to even larger problem-scales could be easily implemented via sequence parallelism (Liu et al., 2023), which allows scaling quadratic self-attention to millions of tokens. Note that Transformer/AB-UPT requires less memory due to constant memory complexity of FlashAttention (Dao et al., 2022), whereas Transolver uses memory heavy torch.einsum operations to create slices. See Appendix A.7 for more details on Transolver scaling.
@@
While such approaches have been explored in the past (Alkin et al., 2024a; b), many neural surrogate architectures focus on linear attention mechanisms (Li et al., 2023a; Wu et al., 2024) or computation parallelization (Luo et al., 2025) to be able to train with millions of inputs. However, such approaches would still be prohibitively expensive. For example, extrapolating the reported GPU requirements for training Transolver++ (Luo et al., 2025) to the 150 million input points from DrivAerML (Ashton et al., 2024b) would necessitate roughly 256 GPUs.
We show in Figure 5 that it is feasible to train standard Transformer self-attention to accurately predict simulation results on million-scale meshes by training on lower resolutions, outperforming linear transformer architectures (such as Transolver). As self-attention requires quadratically more compute w. r. t. the number of inputs, computational demands increase. However, the number of GPU-hours required to train models is still low (models with 16 thousand surface/volume points train in less than 10 GPU-hours). Given even more complicated physical simulations, it would be no issue to scale our AB-UPT to more GPUs via sequence parallelism (Liu et al., 2023), which distributes tokens among multiple GPUs, to allow processing millions of tokens despite the quadratic complexity of self-attention. Due to the low amount of GPU-hours required, even for physical dynamics that arise from meshes with over 100 million cells, AB-UPT could be easily scaled to even more complicated problems. For reference, foundation models in domains such as computer vision (e. g., Oquab et al. (2023)) use 10 to 100 thousand GPU-hours, and compute requirements for language models can be millions of GPU-hours (e. g., Dubey et al. (2024)).
Note that random subsampling, as we apply it to create a low-resolution input mesh, is also a form of data augmentation where a network will never see the same constellation of inputs twice during training. As high-quality numerical CFD simulations are expensive to obtain (e. g., DrivAerML costs \$1M, see Table 2), datasets often have less than a thousand samples, where this form of regularization helps to prevent overfitting.
@@ 4.3 Model design
We design AB-UPT, a successor of UPT that preserves useful properties of UPT, such as using a neural field for decoding and being mesh independent (i.e., the input is decoupled from the output). We observe (in Table 3) that UPT is not able to model complicated dynamics such as the vorticity well, and gradually change the UPT architecture to employ techniques that allow efficient scaling to large problem sizes and obtain significantly higher accuracies than a vanilla UPT. We visualize design choices in Table 3 using AhmedML and motivate the design decisions below. Appendix A.2 validates this study on DrivAerML. Additional micro-design choices for the geometry branch are presented in Appendix A.8.
Table 3: Impact of design steps for transforming UPT into AB-UPT. Performance is measured as relative L2 error (in %) for surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$ and volume vorticity $\bm{\omega}$. Train speed measures seconds per 100 update steps with batch size 1. Test speed measures seconds to create a prediction for the full simulation mesh. Speeds are measured on a NVIDIA H100 GPU. The study is conducted on AhmedML.
@@
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
baseline UPT 4.38 3.20 37.13 7.0 0.2
macro + large decoder 4.25 2.73 15.03 14.7 1.4
design + decoder-only with self-attn 3.41 2.09 6.76 14.9 163
re-add + anchor attention 3.41 2.09 6.76 14.9 2.6
properties + decouple geometry encoding 3.31 2.09 6.64 16.1 2.5
micro + split surface/volume branch 3.50 2.02 6.91 11.0 1.5
design + cross-branch interactions 3.35 2.08 6.76 11.0 1.5
+ branch-specific decoders 3.01 1.90 6.52 11.0 1.5
@@
Starting with a UPT model, we train it on AhmedML (Ashton et al., 2024a) and evaluate surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$, and volume vorticity $\bm{\omega}$. UPT struggles with complex decoding tasks (e. g., decoding vorticity) as the decoder consists of a single cross-attention block, and hence, the vast majority of the compute is allocated to encoding.
@@
Subsequently, we aim to improve speed and accuracy. For automotive CFD simulations, we aim to predict surface and volume variables where we use an equal amount of surface and volume tokens. However, these two modalities differ in the way they are calculated in a classical simulation where surface variables have boundary condition $\bm{u}=0$. Therefore, we split surface and volume positions into separate branches (“+ split surface/volume branch”). This halves the number of tokens of each individual branch, resulting in faster training/inference speed and decreased performance, particularly for surface variables.
@@ 4.4 Benchmarking AB-UPT against other neural surrogate models
Table 4: Relative L2 errors (in %) of surface pressure $\bm{p}_{s}$, volume velocity $\bm{u}$ and volume vorticity $\bm{\omega}$. For each model, we indicate if the model has the neural field property and whether they are mesh independent. Lower percentage values indicate better performance in terms of L2 error. We provide the results for ShapeNet-Car, AhmedML, and DrivAerML. AB-UPT outperforms other neural surrogate models, often by quite a margin.
ShapeNet-Car AhmedML DrivAerML Neural field Mesh independent
$\bm{p}_{s}$ $\bm{u}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
PointNet 12.01 3.05 8.51 5.43 67.29 23.26 28.12 880.0
GRAPH U-NET 10.33 2.48 6.42 4.14 53.53 16.13 17.98 330.0
GINO 13.27 2.52 7.90 6.23 71.80 13.03 40.58 98.5
LNO 9.05 2.29 12.91 7.56 72.04 20.51 23.27 173.0
UPT 6.41 1.49 4.25 2.73 15.03 7.44 8.74 90.2
OFormer 7.05 1.60 3.55 2.22 7.78 4.48 6.64 60.4
Transolver 6.46 1.62 3.45 2.05 8.22 4.81 6.78 38.4
Transformer 4.86 1.17 3.41 2.09 6.76 4.35 6.21 47.9
AB-UPT 4.81 1.16 3.01 1.90 6.52 3.82 5.93 35.1
@@
In this section, we compare the performance of AB-UPT against established neural-surrogate models: Table 4 presents the relative L2 error for surface pressure, volume velocity, and vorticity. We compare AB-UPT against the following models: PointNet (Qi et al., 2017a), Graph U-Net (Gao & Ji, 2019), GINO (Li et al., 2023b), LNO (Wang & Wang, 2024), UPT (Alkin et al., 2024a), OFormer (Li et al., 2023a), Transolver (Wu et al., 2024), Transformer (Vaswani et al., 2017). We do not consider Transolver++ as a baseline due to reproducibility issues, which we explain in Appendix B.4 and C.2. These comparisons are conducted across three diverse datasets: ShapeNet-Car (Umetani & Bickel, 2018), AhmedML (Ashton et al., 2024a), and DrivAerML (Ashton et al., 2024b).
Experimental setup
For ShapeNet-Car, we use the same dataset split as in (Wu et al., 2024), with 789 simulation samples for training and 100 for testing; for AhmedML and DrivAerML, we use a random split of 400 training and 50 test samples, and the remaining samples (50 for AhmedML and 34 for DrivAerML) are used for validation. For a full comparison against baselines, we use 4K/4K, 16K/16K, 16K/16K surface/volume points for ShapeNet-Car, AhmedML, and DrivAerML, respectively. When computing metrics, we generate predictions using this number of inputs and repeat the process as many times until the full point cloud is processed. We then compute the evaluation metrics on the resulting predictions.
The 16K/16K surface/volume points for AhmedML and DrivAerML are motivated by the insights of Figure 5, showing that 16K obtains good results at relatively low compute costs. Evaluation procedure and evaluation metrics are detailed in Appendix B. Appendix B also presents the training hyperparameters and the number of model parameters for each model. Additionally, Appendix C.1 provides a statement on the fairness of our experimental setup in comparison to (Wu et al., 2024). Comprehensive results, including wall shear stress and volume pressure, are reported in Appendix A.4. Additional results for varying the number of slices in Transolver is provided in Appendix A.7 and comparison against Erwin (Zhdanov et al., 2025) is conducted in Appendix A.5.
Performance comparison
@@
Table 4 shows that AB-UPT consistently outperforms all other neural surrogate models across every evaluated metric and dataset, often by a substantial margin. Moreover, we observe that most baselines struggle with properly modeling the complex vorticity. AB-UPT, however, maintains strong performance for both surface and volume metrics, highlighting its robustness. For most metrics, either a plain Transformer or Transolver is the second-best-performing model. However, due to the full self-attention mechanism of the Transformer, this model will not be able to scale to industry-standard simulation meshes (Transformer corresponds to the third row of Table 3, which shows roughly 100 times slower test speed than AB-UPT). Transolver, on the other hand, would require sequence parallelism over multiple GPUs to handle large simulation meshes (as has been explored in (Luo et al., 2025)). We can conclude that AB-UPT is the best performing model for neural-surrogate automotive CFD simulation, while also having favorable properties concerning problem scaling.